# Integrating Biological and Machine Intelligence: Attention Mechanisms in Brain-Computer Interfaces

### v1

@@ 6 Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62276169, in part by the Medical-Engineering Interdisciplinary Research Foundation of Shenzhen University under Grant 2024YG008, in part by the Shenzhen University-Lingnan University Joint Research Programme, in part by the Shenzhen-Hong Kong Institute of Brain Science-Shenzhen Fundamental Research Institutions under Grant 2023SHIBS0003, in part by the STI 2030-Major Projects 2021ZD0200500, and in part by the Open Research Fund of the State Key Laboratory of Brain-Machine Intelligence, Zhejiang University (Grant No. BMI2400008).

### v2

@@ 6 Acknowledgments
This work was supported by the National Natural Science Foundation of China (62276169), Medical-Engineering Interdisciplinary Research Foundation of Shenzhen University (2023YG004), Shenzhen University Lingnan University Joint Research Programme, Shenzhen-Hong Kong Institute of Brain Science-Shenzhen Fundamental Research Institutions (2023SHIBS0003), the STI 2030-Major Projects (2021ZD0200500), the Open Research Fund of the State Key Laboratory of Brain-Machine Intelligence, Zhejiang University (Grant No. BMI2400008), and the Shenzhen Science and Technology Program (No. JCYJ20241202124222027 and JCYJ20241202124209011).